find the width of rectangle length is 23.3m and perimeter is 96.8
Found 2 solutions by Alan3354, checkley77:
Answer by Alan3354(69443)
Answer by checkley77(12844)

P=2L+2W
96.8=2*23.3+2W
96.8=46.6+2W
2W=96.8-46.6
2W=50.2
W=50.2/2
W=25.1 ANS. FOR THE WIDTH
PROOF:
96.8=2*23.3+2*25.1
96.8=46.6+50.2
96.8=96.8